OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

rishavraj (rishavraj):

2x + 16 = 12 find x......

OpenStudy (anonymous):

set your function equal to 12 and then solve for x: \[2(x+8) =12\] in order to get rid of the parentheses distribute the 2 to the x and 8 \[2x+16=12\] and solve for x.

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yup that's correct, you can check by just plugging in that value into your original function and seeing if it comes out to what you expect it to (12). \[f(-2) = 2(-2+8)\] \[f(-2) = 2(6)=12\]